# Piezoelectric actuation for integrated photonics

### v1

@@ 6.3 Integrated Acousto-Optic Frequency Shifter
where (), (), and () are the momentum, angular frequency, and velocity of acoustic (optical) wave, respectively. The intensity of the diffracted light can be controlled by modulating the diffraction efficiency which is related with the power and frequency of the RF drive [404]. This capability is conventionally utilized in the Q-switch laser [405] and in the generation of optical pulses with ultra-high on-off contrast [355]. The frequency shifting of the acoustically scattered light naturally leads to the desirable single sideband suppressed carrier (SSB-SC) modulation, which is complicated to implement using electro-optical modulators [406]. This unique feature of AOM has been translated into an indispensable optical device, acousto-optic frequency shifter (AOFS), which is widely used in the heterodyne detection [134, 407].
@@
Despite of the ubiquitous application of AOM, current commercially available AOM is fabricated on a stand-alone bulk piezoelectric substrate and the acoustic frequency is limited to a few hundreds of megahertz. The required RF drive power is at Watts scale due to the large acoustic and optical mode volume. The implementation of Bragg diffraction on integrated planar photonics has been demonstrated on a released AlN membrane for the first time [408]. Despite the low scattering efficiency, asymmetric modulation with 11 GHz frequency shift is achieved with only 100 mW RF power. Recently, the efficiency is improved by exciting SAW on an unreleased LiNbO3 thin film, which has higher piezoelectric response and power handling capability [137, 404], as shown in Fig. 49. The input and output light is guided through either an etched LiNbO3 rib waveguide [137] or a bound-in-continuum (BIC) waveguide mode formed by patterned polymer on LiNbO3 film [404]. The light from the waveguide transits into a two-dimensional Gaussian beam confined in the slab formed by the LiNbO3 thin film, which interacts with the SAW in the same region. An ideal AOFS would require most of the input light to be shifted to the desired frequency with zero power remaining in the original frequency or scattered to other sidebands. This behavior can be characterized by the modulation efficiency (the fraction of input light that is scattered to the target sideband), and carrier and sideband suppression ratios. The optical spectrum of the shifted light is shown in Fig. 49(b, d). It can be found that most optical power is in the anti-Stokes sideband with around 3 GHz higher frequency, and the center carrier and other sideband is 30 dB and 40 dB smaller, respectively. The efficiency is measured to be around 3 per 1 W RF drive. The frequency shifting to the Stokes sideband can be realized by changing to the other input port [137, 404]. The optical bandwidth is proportional to , where , , and are the optical wavelength, period of IDT fingers, and width of the IDT, respectively [137]. Optical bandwidths of 14 nm and 35 nm are achieved for the works in Ref. [137] and [404], respectively. These integrated AOFS have compact size, low power consumption, and gigahertz operation, which can be used for on-chip heterodyne detection, microwave signal processing, and optical signal routing.
@@
For the same device in Fig. 46(e), where a pair of TE and TM modes of a Si3N4 microring resonator are coupled by HBAR, AOFS can be realized by designing the optical resonator in overcoupled regime. The overcoupling is required to have high extraction efficiency of the frequency shifted light. The strong coupling between the two modes leads to complete optical conversion when cooperativity equals one. The optical transmission when injecting TE light into the waveguide and under 20 dBm RF drive is shown in the upper panel of Fig. 50(a). At the optical resonance, 50% of the input light is converted into TM mode, which has a higher frequency. Since the device is not in the strong overcoupled regime, the rest 50% is dissipated as the intrinsic loss of the optical resonator. The conversion is also bidirectional, where the input TM mode can be converted into TE mode (see the lower panel of Fig. 50(a)). The asymmetry of the modulation sidebands can be seen from the output optical spectrum as shown in Fig. 50(b). A single frequency laser at TE mode is input and modulated with a fixed frequency microwave drive. The spectrum is broad because of the limited resolution of the spectrum analyzer. The main peak of the output TM light shifts to a lower wavelength. The light remaining in the carrier and other sideband is more than 10 dB smaller. The frequency shift can be seen more clearly on the linear scale. The choice of TE and TM mode pair rotates the polarization in addition to the frequency shift, which also makes it convenient to separate the converted light from the carrier.
@@
In addition to the microring resonator, AOFS is also performed on a straight Si waveguide, the same device shown in Fig. 47(c). A pair of symmetric and anti-symmetric waveguide modes is coupled by a SAW excited from the AlN IDT. As can be seen in Fig. 50(c), a Rabi-like oscillation between the two modes is observed while the RF drive power is swept, demonstrating complete mode conversion. The optical spectrum of the output light at different frequencies is illustrated in Fig. 50(d). The Stokes sideband dominates and is more than 30 dB larger than the carrier and anti-Stokes sideband. The Si waveguide supports a broad optical bandwidth of 1.7 nm, in stark contrast to the microring resonator. The RF bandwidth is limited by the bandwidth of the IDT to a few MHz (see Fig. 50(e)), which can be improved in the future by adopting chirped IDT design [409].
@@
While the integrated optical waveguides help to extend the interaction length between photon and phonon on chip, the confining of acoustic wave could further improve the propagation length and interaction strength. In that sense, the development of acoustic waveguide and phononic circuits has drawn much attention recently, and major progresses have been made in both released [324] and unreleased [302, 301] waveguiding. Furthermore, the co-propagation of phonon and photon in the same waveguide has been demonstrated successfully, paving the path for furture optomechanical circuits [410, 411].
@@
Apart from the classical application in heterodyne detection, AOFS will find application in future frequency-domain optical quantum information processing and quantum computing based on frequency-bin optical qubits [412, 413]. As demonstrated in Fig. 50(c), the conversion between the two optical modes can be controlled by the RF power such that the output power is equally split into the two modes. Thus, a frequency domain beam splitter can be built which is the building block for optical quantum computing [414, 415]. The frequency beam splitter is conventionally constructed by cascading electro-optical modulators and optical pulse shaper [416, 417], which is bulky and power consuming, making it difficulty for scaling. Therefore, it can be foreseen that the integrated AOFS will become an indispensable component for future large scale optical quantum computing. So far, the experiments of AOFS are done in the classical domain, it will be necessary to explore the quantum behavior of it, such as the operation with non-classical light, quantum coherence and interference [418], and quantum fidelity [416].
@@ 7.1 Hybrid integrated quantum systems
Hybrid integrated quantum systems have attracted extensive efforts recently, as they take advantages of distinct quantum objects for high-performance quantum applications [419, 420]. For example, the superconducting (SC) qubit has become one of the mature platforms for quantum computing. However, the operation at cryogenic temperature limits its scalability. On the other hand, optical photon is good at quantum information communication with low quantum noise, which makes it possible for scalable quantum network [372]. In contrast to the electromagnetic wave, the low speed of acoustic wave endows it with sub-micron wavelength at microwave frequencies, which allows compact device size. Moreover, the long coherence time (low losses) of the mechanical resonator is suitable to store quantum information. A hardware-efficient Quantum Random Access Memory (QRAM) has been theoretically proposed [421]. Therefore, hybrid quantum system built from these different modules paves the way for a multi-functional and scalable quantum processor. However, achieving the transduction between different domains with high quantum efficiency and low added noise is still challenging.
@@ 7.2 Piezoelectric programmable photonic circuits
The key component of the programmable MZI is the optical phase shifter. For better performance of the MZI, an ideal phase shifter would require a high-speed response (large bandwidth), low power consumption, and compact size. The latter two metrics together determine the scalability of the Mach-Zehnder mesh (MZM). To achieve the unitary transformation of input channels, total of MZI will be needed [231]. This scaling places a restrict requirement on the power consumption and size of the designed phase shifter. Conventionally, the MZI is thermally tuned to build the programmable circuit, from Si ONN [55] to the implementation of quantum algorithms on a Si3N4 photonic circuits [56]. Although the thermal phase shifter is easy to design and fabricate, its actuation speed is slow (millisecond) and power consumption is large (milliwatt). Moreover, the thermal tuning is incompatible with cryogenic experiments, such as co-integration with SC circuits, single photon detector, or trapped ion. Modulation of the Si waveguide through free carrier injection has been demonstrated with low power and high speed [427]. However, the light propagation loss is too high for large-scale integration. Although the electrostatic phase shifter provides a solution for low power and compact size, its bandwidth is limited by the mechanical resonance to 10-100 kHz [100, 87, 104, 428]. Therefore, a lower power, low optical loss, high-speed, and compact-size phase shifter will become essential for future programmable photonic circuits.
@@ 7.3 Synthetic frequency dimensions in photonics
In addition to studying the topological effects, the synthetic frequency dimension has been utilized to implement an arbitrary linear transformation by engineering the coupling between frequency modes [451]. Similar to the MZM presented in the last section, the input optical states can be arbitrarily mapped to the output but in the frequency space. The advantages of frequency domain implementation include the compact device size and scalability, since all optical channels transmit in the same optical waveguide. Recently, matrix vector multiplication has been demonstrated by acousto-optic modulating a single Si photonic crystal resonator [446], as shown in Fig. 54(d, e). In the future, linear optical quantum information processing in the frequency dimension can be achieved by incorporating non-classical light. Although experimental realizations using bulk EOM and pulse shaper have been achieved [417, 416], chip-scale photonic circuit and AOM could enable large-scale optical computing with high-fidelity frequency-bin logic gates [451, 446]. Additionally, a non-unitary transformation can be realized by introducing the modulation of gain or loss of the optical resonator, which is generally the case in neural networks [451]. Finally, combining with the optical nonlinearity of the photonic waveguide, such as Si3N, Si, and LiNbO, novel dynamics of the soliton generation has been theoretically predicted [452]. The nonlinearity is also required for implementing deep learning algorithms, a step for fully optical neural networks [55]. Therefore, we believe the integrated AOM will play an important role in both classical and quantum optical computing in the synthetic frequency dimension through efficient modulation and frequency conversion.

### v2

@@ 6.3 Integrated Acousto-Optic Frequency Shifter
where (), (), and () are the momentum, angular frequency, and velocity of acoustic (optical) wave, respectively. The intensity of the diffracted light can be modulated by controlling the diffraction efficiency which is related with the power and frequency of the RF drive [423]. This capability is conventionally utilized in the Q-switch laser [424] and in the generation of optical pulses with ultra-high on-off contrast [371]. The frequency shifting of the acoustically scattered light naturally leads to the desirable single sideband suppressed carrier (SSB-SC) modulation, which is complicated to implement using electro-optical modulators [425]. This unique feature of AOM has been translated into an indispensable optical device, acousto-optic frequency shifter (AOFS), which is widely used in the heterodyne detection [143, 426].
@@
Despite of the ubiquitous application of AOM, current commercially available AOMs are fabricated on a stand-alone bulk piezoelectric substrate and the acoustic frequency is limited to a few hundreds of megahertz. The required RF drive power is at Watts scale due to the large acoustic and optical mode volume. The implementation of Bragg diffraction on integrated planar photonics has been demonstrated on a released AlN membrane for the first time [427]. Despite the low scattering efficiency, asymmetric modulation with 11 GHz frequency shift is achieved with only 100 mW RF power. Recently, the efficiency is improved by exciting SAW on an unreleased LiNbO3 thin film, which has higher piezoelectric response and power handling capability [146, 423], as shown in Fig. 49. The input and output light is guided through either an etched LiNbO3 rib waveguide [146] or a BIC waveguide mode formed by patterned polymer on LiNbO3 film [423]. The light from the waveguide transits into a two-dimensional Gaussian beam confined in the slab formed by the LiNbO3 thin film, which interacts with the SAW in the same region. An ideal AOFS would require most of the input light to be shifted to the desired frequency with zero power remaining in the original frequency or scattered to other sidebands. This behavior can be characterized by the modulation efficiency (the fraction of input light that is scattered to the target sideband), and carrier and sideband suppression ratios. The optical spectrum of the shifted light is shown in Fig. 49(b, d). It can be found that most optical power is in the anti-Stokes sideband with around 3 GHz higher frequency, and the center carrier and the Stokes sideband is 30 dB and 40 dB smaller, respectively. The efficiency is measured to be around 3 per 1 W RF drive. Frequency shifting to the Stokes sideband can be realized by changing the input port [146, 423]. The optical bandwidth is proportional to , where , , and are the optical wavelength, period of IDT fingers, and width of the IDT, respectively [146]. Optical bandwidths of 14 nm and 35 nm are achieved for the works in Ref. [146] and [423], respectively. These integrated AOFS have compact size, low power consumption, and gigahertz operation, which can be used for on-chip heterodyne detection, microwave signal processing, and optical signal routing.
@@
For the same device in Fig. 46(e), where a pair of TE00 and TM00 modes of a Si3N4 microring resonator are coupled by HBAR, AOFS can be realized by designing the optical resonator in overcoupled regime. The overcoupling is required to have high extraction efficiency of the frequency shifted light. The strong coupling between the two modes leads to complete optical conversion when cooperativity equals one. The optical transmission when injecting TE00 light into the waveguide and under 20 dBm RF drive is shown in the upper panel of Fig. 50(a). At the optical resonance, 50% of the input light is converted into TM00 mode, which has a higher frequency. Since the device is not in the strong overcoupled regime, the rest 50% is dissipated as the intrinsic loss of the optical resonator. The conversion is also bidirectional, where the input TM00 mode can be converted into TE00 mode (see the lower panel of Fig. 50(a)). The asymmetry of the modulation sidebands can be seen from the output optical spectrum as shown in Fig. 50(b). A single frequency laser at TE00 mode is input and modulated with a fixed frequency microwave drive. The spectrum is broad because of the limited resolution of the spectrum analyzer. The main peak of the output TM00 light shifts to a lower wavelength. The light remaining in the carrier and other sideband is more than 10 dB smaller. The frequency shift can be seen more clearly on the linear scale. The choice of TE00 and TM00 mode pair rotates the polarization in addition to the frequency shift, which also makes it convenient to separate the converted light from the carrier using commercial polarization beam-splitter.
@@
In addition to the microring resonator, AOFS is also performed on a straight Si waveguide, the same device shown in Fig. 47(c). A pair of symmetric and anti-symmetric waveguide modes is coupled by a SAW excited from the AlN IDT. As can be seen in Fig. 50(c), a Rabi-like oscillation between the two modes is observed by sweeping the RF drive power, demonstrating complete mode conversion. The optical spectrum of the output light at different sidebands is illustrated in Fig. 50(d). The Stokes sideband dominates and is more than 30 dB larger than the carrier and anti-Stokes sideband. The Si waveguide supports a broad optical bandwidth of 1.7 nm, in stark contrast to the microring resonator. The RF bandwidth is limited by the bandwidth of the IDT to a few MHz (see Fig. 50(e)), which can be improved in the future by adopting chirped IDT design [428].
@@
While the integrated optical waveguides help to extend the interaction length between photon and phonon on chip, the confining of acoustic wave could further improve the propagation length and interaction strength. In that sense, the development of acoustic waveguide and phononic circuits has drawn much attention recently, and major progress has been made in both released [340] and unreleased [318, 317] waveguiding. Furthermore, the co-propagation of phonon and photon in the same waveguide has been demonstrated successfully, paving the path for furture optomechanical circuits [429, 430].
@@
Apart from the classical application in heterodyne detection, AOFS will find application in future frequency-domain optical quantum information processing and quantum computing based on frequency-bin optical qubits [431, 432]. As demonstrated in Fig. 50(c), the conversion between the two optical modes can be controlled by the RF power such that the output power is equally split into the two modes. Thus, a frequency domain beam splitter can be built which is the building block for optical quantum computing [433, 434]. The frequency beam splitter is conventionally constructed by cascading electro-optical modulators and optical pulse shaper [435, 436], which is bulky and power consuming, making it difficult to scale. Therefore, it can be foreseen that the integrated AOFS will become an indispensable component for future large-scale optical quantum computing. So far, the experiments of the aforementioned AOFS are done in the classical domain, it will be necessary to explore the quantum behavior of it, such as the operation with non-classical light, quantum coherence and interference [437], and characterization of the quantum fidelity [435].
Besides the frequency shifts caused by Brillouin scattering between optical and acoustic modes, frequency shifts can also occur through linear phase modulation of the optical carrier , resulting in a photon frequency shift by . This phenomenon is a key aspect of the Fourier transform relationship between frequency and time, commonly referred to as spectral shearing in the literature [439]. Optical frequency shifters utilizing spectral shearing have been demonstrated with both bulk and integrated electro-optic modulators [439, 440] as well as piezo-optomechanical phase modulators [438, 441]. Figure 51 (a, b) illustrate an optomechanical frequency shifter, where a suspended AlN waveguide is stretched piezoelectrically which then changes the effective refractive index of the waveguide and the phase experienced by the photon [438]. The mechanical thickness mode of the AlN waveguide is activated, enhancing modulation efficiency through mechanical resonance. In most cases, the phase is modulated sinusoidally at RF frequencies. To achieve linear phase modulation, it is essential that optical pulses are significantly shorter than the sinusoidal wave period, ensuring that the phase during the pulse is linearly modulated. Consequently, this approach is not suitable for long optical pulses or CW light, as it would lead to spectral distortion and the generation of multiple sidebands [438]. It is noteworthy that the frequency shift does not depend directly on the modulation frequency itself, but on the rate of change of the phase over time, which is influenced by the modulation depth, modulation frequency, and the relative phase between the RF drive and the photon arrival time. The dependence of the frequency shift on the RF drive phase is experimentally measured, as shown in Fig. 51(c). Over 300 GHz frequency shift has been achieved with near-unity intrinsic efficiency under 2 W RF drive [438]. The preservation of quantum coherence after frequency shifting is further demonstrated by conducting the Hong-Ou-Mandel interference between two photons with different colors, as shown in Fig. 51(d). In the experiment, one red photon is blue-shifted to have the same frequency as the other. The large dip of the coincidence rate between two single photon detectors indicates indistinguishability of the two photons after the frequency shift (blue curve in Fig. 51(d)). If the red photon is unshifted (black) or red-shifted (red), the visibility becomes low. This optomechanical frequency shifter is broadband and does not require pump light, which makes it readily be applicable in optical quantum communication and information processing. In the future, the study of new mechanical modes (e.g., SAW and BAW), photonic crystal waveguide, and optical materials with larger photoelastic coefficients could further improve the modulation efficiency and reduce the device footprint by stronger optomechanical interaction strength.
@@ 7.1 Hybrid Integrated Quantum Systems
Hybrid integrated quantum systems have attracted extensive efforts recently, as they take advantages of distinct quantum objects for high-performance quantum applications [442, 443]. For example, the SC qubit has become one of the mature platforms for quantum computing. However, the operation at cryogenic temperature limits its scalability. On the other hand, optical photon is good at quantum information communication with low quantum noise, which makes it possible for scalable quantum network [390]. In contrast to the electromagnetic wave, the low speed of acoustic wave endows it with sub-micron wavelength at microwave frequencies, which allows compact device size. Moreover, the long coherence time (low losses) of the mechanical resonator makes it suitable to store quantum information. A hardware-efficient Quantum Random Access Memory (QRAM) has been theoretically proposed [444]. Therefore, hybrid quantum system built from these different modules paves the way for a multi-functional and scalable quantum processor. However, achieving the transduction between different domains with high quantum efficiency and low added noise is still challenging.
@@ 7.2 Piezoelectric Programmable Photonic Circuits
The key component of the programmable MZI is the optical phase shifter. For better performance of the MZI, an ideal phase shifter would require a high-speed response (large bandwidth), low power consumption, and compact size. The latter two metrics together determine the scalability of the MZM. To achieve the unitary transformation of input channels, total of MZI will be needed [243]. This scaling places a restrict requirement on the power consumption and size of the designed phase shifter. Conventionally, the MZI is thermally tuned to build the programmable circuit, from Si ONN [57] to the implementation of quantum algorithms on a Si3N4 photonic circuits [58]. Although the thermal phase shifter is easy to design and fabricate, its actuation speed is slow (millisecond) and power consumption is large (milliwatt). Moreover, the thermal tuning is incompatible with cryogenic experiments, such as co-integration with SC circuits, single photon detector, or trapped ion. Modulation of the Si waveguide through free carrier injection has been demonstrated with low power and high speed [450]. However, the light propagation loss is too high for large-scale integration. Although the electrostatic phase shifter provides a solution for low power and compact size, its bandwidth is limited by the mechanical resonance to 10-100 kHz [102, 89, 106, 451]. Therefore, a lower power, low optical loss, high-speed, and compact-size phase shifter will become essential for future programmable photonic circuits.
@@ 7.3 Synthetic Frequency Dimensions in Photonics
In addition to studying the topological effects, the synthetic frequency dimension has been utilized to implement an arbitrary linear transformation by engineering the coupling between frequency modes [474]. Similar to the MZM presented in the last section, the input optical states can be arbitrarily mapped to the output but in the frequency space. The advantages of frequency domain implementation include the compact device size and scalability, since all optical channels transmit in the same optical waveguide. Recently, matrix vector multiplication has been demonstrated by acousto-optic modulating a single Si photonic crystal resonator [469], as shown in Fig. 55(d, e). In the future, linear optical quantum information processing in the frequency dimension can be achieved by incorporating non-classical light. Although experimental realizations using bulk EOM and pulse shaper have been achieved [436, 435], chip-scale photonic circuit and AOM could enable large-scale optical computing with high-fidelity frequency-bin logic gates [474, 469]. Additionally, a non-unitary transformation can be realized by introducing the modulation of gain or loss of the optical resonator, which is generally the case in neural networks [474]. Finally, combining with the optical nonlinearity of the photonic waveguide, such as Si3N, Si, and LiNbO3, novel dynamics of the soliton generation has been theoretically predicted [475]. The nonlinearity is also required for implementing deep learning algorithms, a step for fully optical neural networks [57]. Therefore, we believe the integrated AOM will play an important role in both classical and quantum optical computing in the synthetic frequency dimension through efficient modulation and frequency conversion.
7.4 Novel Piezoelectric Materials
As a crucial element in piezoelectric actuators, investigating new materials with improved piezoelectric properties will significantly boost the efficiency, reduce the size, and lower the power consumption of these devices. Benefiting from the advancement of high-quality thin film deposition on semiconductor substrates, a number of good candidates have received intensive research efforts in the past decade, such as AlScN [128, 296], BaTiO3 (BTO) [110], hafnium zirconium oxide (HZO) [246, 477], and potassium–sodium niobate (KNN) [478]. Among them, AlScN has witnessed promising progress in recent years, where a variety of thin-film deposition techniques have been explored, from reactive sputter deposition, molecular beam epitaxy (MBE), metal organic chemical vapor deposition (MOCVD), and pulsed laser deposition (PLD). A review of these deposition methods is provided in Ref. [128]. Compared to AlN, dry etching of AlScN is more challenging due to the low volatility of ScCl3, especially for high Sc doping concentration [479]. Nevertheless, inductively coupled plasma reactive ion etching (ICP-RIE) of AlScN has been studied in Ref. [480]. Alternatively, AlScN can be etched physically by Ar+ ion milling [481]. Wet etching of AlScN has also been studied with various wet chemical solutions, such as alkaline solutions tetramethyl ammonium hydroxide (TMAH) and potassium hydroxide (KOH) and acid solution phosphoric acid (H3PO4) [479, 482]. The integration of CMOS-compatible AlScN PICs at the wafer scale has been achieved [483, 484], together with enhanced second-order nonlinearity [485, 486] and electro-optic modulation [487, 488]. Efficient AOMs based on AlScN have been demonstrated with various structure designs [489, 476, 490, 491]. An example of the AOM is shown in Fig. 56(a), where a SAW is excited by the IDT fabricated on the AlScN thin film [476]. A Si waveguide is modulated by the SAW and a voltage-length product of 2.34 Vcm is reported, which is still comparable to that of AlN (see Table 5).
In the past decade, BTO has attracted considerable research interest because of its substantial Pockels and piezoelectric coefficients (see Table 1) and the successful deposition of high-quality thin films on silicon substrates [83, 162]. CMOS-compatible wet etching via buffered hydrogen fluoride (HF) [492] and dry etching through fluorine-based [493] and Ar-based [494] plasma have made major progresses. Further optimization of the etching conditions would allow for photonic and optomechanical circuits made from patterned BTO. Integrated electro-optic modulators made from BTO have achieved over 20 GHz modulation bandwidth in both room [84] and cryogenic temperatures [85], as shown in Fig. 56(b, c). Nonetheless, BTO’s significant piezoelectric response has not been sufficiently investigated for applications in piezoelectric actuation and AOM. Alongside its ferroelectric characteristics, BTO will pave the way for new tunable and reconfigurable electro-optic and acousto-optic devices [86].
Since its first report of ferroelectricity in 2011 [495], HZO has received intensive research interest by serving as ferroelectric transistors and memories. Atomic layer deposition (ALD) of HZO thin films, being compatible with CMOS process, has become a commonly used method [496]. However, the thickness of these films is typically limited to several tens of nanometers. The recently developed atomic layer etching (ALE) technique enables precise control of the etching of HZO thin films [497]. Beyond microelectronic devices, high Q-factor electromechanical resonator based on an ultra-thin HZO film (10 nm) has been demonstrated [498]. Even though it is still in its early stages, significant progress has been made in applying HZO in integrated photonic circuits. For example, a nonvolatile optical phase shifter utilizing a hybrid Si3N4-HZO optical waveguide has been realized, paving the way for low-power programmable photonic circuits in the future [499]. Recently, an ultra-low power piezoelectric tuning of a hybrid Si-HZO optical ring resonator has been demonstrated with 1 GHz/V tuning efficiency and 1 pW/MHz power consumption [246], as illustrated in Fig. 56(d-f). The comparison with AlN and PZT is included in Table 3, where the tuning efficiency of HZO is much larger than that of AlN and comparable to PZT. However, the optical waveguide loss is still high (48 dB/cm), primarily due to the roughness of the HZO film and the surface of amorphous Si film [246].
Due to the demand for eco-friendly, lead-free piezoelectric materials, KNN has emerged as a viable substitute for conventional piezoelectric materials like PZT. Superior piezoelectric response of KNN ceramics with pC/N [500] and KNN single crystal with pC/N [501] have been reported. KNN thin-films are generally synthesized via RF magnetron sputtering [502], sol-gel [503], and chemical solution deposition techniques [504]. A post-annealing process at around 700 ∘C is required for the KNN film to crystallize [502]. The precise regulation of the composition of KNN thin films through optimized processes is responsible for the enhanced piezoelectric response of KNN [505]. Various etching methods have been explored including wet etching with alkaline solution [506] and drying plasma etching [507]. Owing to their piezoelectric characteristics that are on par with PZT, KNN thin films are being investigated for various applications [508], such as sensors, actuators [506], and devices for energy harvesting. Future research on further improving the performance and reliability of KNN films will facilitate its application in photonic circuits for piezoelectric tuning and AOM.